OpenStudy (anonymous):

FAN & MEDAL. In your own words explain the difference between the following two expressions: (-1)^2/4 and (-1)^1/2

OpenStudy (solomonzelman):

I have a question. \[\frac{2}{4}~~~~~~~~=~or~≠~~~(t o)~~~~~~~~\frac{1}{2}~~~~?\]

OpenStudy (anonymous):

2/4 = 1/2

OpenStudy (anonymous):

@SolomonZelman